What is the velocity of a 0.5 kg ball that has a momentum of 3.0 kg m/s?


1
Expert's answer
2022-04-27T13:33:43-0400
"v=\frac{p}{m}=\rm\frac{3.0\: kg\cdot m/s}{0.5\: kg}=6.0\: m/s"